OpenStudy (anonymous):

Who can rationalize g^3/4 h^5/4 over square root of 3r^3/4 ?

OpenStudy (anonymous):

mutiply both the numerator and the denominator with the denominator

Parth (parthkohli):

\[\LARGE {{g^{1 \over 4} * g^{1 \over 4} * g^{1 \over 4} * h * h^{1 \over 4}} \over 3r^{3 \over 2} / 4} \]

OpenStudy (anonymous):

@ParthKohli . Eh? Elaborate pls.

Parth (parthkohli):

You can see my tutorial, you'll understand everything.. If you don't then turn over to me.

OpenStudy (anonymous):

\[{g^{{3}\over{4}}\times h^{{5}\over{4}}}\over{\sqrt{3r^{{3}\over{4}}}}\] \[={{g^{{3}\over{4}}\times h^{{5}\over{4}}}\over\sqrt{3}r^{{3}\over{8}}}\] \[={{g^{{3}\over{4}}\times h^{{5}\over{4}}}\times r^{{3}\over{8}}\over\sqrt{3}}\]

OpenStudy (anonymous):

I don't understand. What the procedure? Like the steps?

Parth (parthkohli):

@sheg square root of 3 is not rationalized. Simplify it further..

Parth (parthkohli):

Ask sheg, he'll tell you.

OpenStudy (anonymous):

How can the radical be removed?

OpenStudy (anonymous):

@sheg ?

Parth (parthkohli):

Multiply both the numerator and denominator by square root of 3.

OpenStudy (anonymous):

\[{g^{{3}\over{4}}\times h^{{5}\over{4}}}\over{\sqrt{3r^{{3}\over{4}}}}\] \[={{g^{{3}\over{4}}\times h^{{5}\over{4}}}\over\sqrt{3}r^{{3}\over{8}}}\] \[={{g^{{3}\over{4}}\times h^{{5}\over{4}}}\times r^{{3}\over{8}}\over\sqrt{3}}\] \[={3^{{-1}\over{2}}}\times {{g^{{3}\over{4}}\times h^{{5}\over{4}}}\times r^{{3}\over{8}}}\]

OpenStudy (anonymous):

@sheg Is it like the final answer?

OpenStudy (anonymous):

on simplification you will get 3^(-1/2)g^(3/4)h^(5/4)r^(-3/8) hey in above i forgot to put minus sign in the power of r @chol

OpenStudy (anonymous):

Ok. Thanks so much @sheg

OpenStudy (anonymous):

did u understood how to solve this one

OpenStudy (anonymous):

Ahm. I kind of understand it now :)

OpenStudy (anonymous):

see sqrt of 3 can be written as\[\sqrt{3} = 3^{{1}\over{2}}\]

OpenStudy (anonymous):

Yeah. I know that thing.

OpenStudy (anonymous):

ok so in case of sqrt(r^(3/4)) = r^(3/8) and as it is in the denominator so it can be written in the numerator as r^(-3/8)

OpenStudy (anonymous):

Oh. That's kind of helpful. Thank you.

OpenStudy (anonymous):

@sheg :)